a clock has its sec hand length 2m.find ave.speed&moduls of its tip in15 sec


 

Grade:12

3 Answers

Akash Kumar Dutta
11 years ago

Dear Sai Shivani,

ave. speed=total dist./total time
                                                   =  pie.r/2.15 = pie.r/30 = pie/15 m/s(ANS)

Regards.                              

muhammed . nabeel
11 years ago

6 degreeee

Prajwal kr
11 years ago

Average speed= (0.5 x pi x radius)/time

                     =0.209 m/s